OpenStudy (anonymous):

Please help !!! Solve. x2 + 15x = -56 x = 8 and x = 7 x = -8 and x = -7 x = -8 and x = 7 x = 8 and x = -7

OpenStudy (paki):

x2+7x+8x+56=0 x(x+7) + 8(x+7)=0 (x+8)(x+7)=0 x+8=0 and x+7=0 x=-8 and x=-7

OpenStudy (anonymous):

Thanks so much, now i'm starting to remember this from algebra 1